# 4DHands: Reconstructing Interactive Hands in 4D with Transformers

### v2

@@ Abstract
In this paper, we introduce 4DHands, a robust approach to recovering interactive hand meshes and their relative movement from monocular inputs. Our approach addresses two major limitations of previous methods: lacking a unified solution for handling various hand image inputs and neglecting the positional relationship of two hands within images. To overcome these challenges, we develop a transformer-based architecture with novel tokenization and feature fusion strategies. Specifically, we propose a Relation-aware Two-Hand Tokenization (RAT) method to embed positional relation information into the hand tokens. In this way, our network can handle both single-hand and two-hand inputs and explicitly leverage relative hand positions, facilitating the reconstruction of intricate hand interactions in real-world scenarios. As such tokenization indicates the relative relationship of two hands, it also supports more effective feature fusion. To this end, we further develop a Spatio-temporal Interaction Reasoning (SIR) module to fuse hand tokens in 4D with attention and decode them into 3D hand meshes and relative temporal movements. The efficacy of our approach is validated on several benchmark datasets. The results on in-the-wild videos and real-world scenarios demonstrate the superior performances of our approach for interactive hand reconstruction. More video results can be found on the project page: https://4dhands.github.io.
Figure 1: The proposed method, 4DHands, can robustly recover interactive hand meshes and their relative movement from monocular inputs.
@@ 1 Introduction
Hand mesh recovery from monocular inputs is a fundamental technology for numerous advanced applications like virtual/augmented reality, human-computer interaction, robotics, and embodied AI. Despite the remarkable progress achieved by existing methods [27, 29, 53] in this field, challenges remain in addressing generalization under in-the-wild scenarios, capturing robust hand motions with temporal stability over time, and recovering plausible interactions between two hands.
With the developments in transformer architectures, recent state-of-the-art approaches [39] to monocular human/hand mesh recovery have shifted toward investigating the usage of simple and high-capacity models along with large-scale in-the-wild data. Despite the improvement in 2D alignment, existing approaches still can not robustly handle close interaction and reconstruct hand meshes in the 4D space with stable temporal movement. Moreover, existing approaches typically handle two-hand and single-hand scenarios differently. As the two-hand status frequently switches between close interaction and separation in real-world scenes, the inconsistent processing procedure inevitably leads to clear jittering artifacts and temporal instability.
In this paper, we identify the three major limitations in previous solutions: i) the lack of a unified solution to handle hand image inputs in various situations, ii) neglecting the positional relationship between the hands within images, and iii) unstable estimation of the temporal translation of hands especially during the status shift between interaction and separation. To address these issues, we proposed 4DHands, a robust solution to recover interactive hand meshes and their relative movements in 4D from monocular inputs. The key insight of our solution is to unify the feature learning of two hands in various scenarios and fully leverage their positional relationship. To achieve this, 4DHands adopts a transformer-based architecture and includes novel tokenization and 4D feature fusion strategies.
@@
Specifically, we first propose a Relation-aware Two-Hand Tokenization (RAT) method to embed positional relation information of each hand into its tokens. Such a tokenization method brings the following benefits threefold. i) RAT is not reliant on the co-occurrence of two hands and can thus handle both single-hand and two-hand inputs in a unified manner. ii) RAT takes the overlapped status of two hands into account and adjusts the tokens accordingly, which allows the network to learn a more concentrated distribution of the hand tokens regardless of the various scales and positions. iii) RAT explicitly encodes the relative positions of two hands into the tokens, facilitating the reconstruction of close and intricate interaction in real-world scenarios. After obtaining the image tokens, we further develop a Spatio-temporal Interaction Reasoning (SIR) module to infer hand poses and the relative positions between the two hands. Thanks to the proposed tokenization method RAT, SIR can fuse both the spatial and temporal information more effectively as the hand tokens indicate the relative relationship of two hands, thereby enhancing the robustness of hand mesh and temporal movement recovery in the 4D space.
@@
  • We propose a Spatio-temporal Interaction Reasoning (SIR) module to fuse 4D hand features. The collaboration of RAT and SIR further enhances the stability of hand mesh and temporal movement recovery in the 4D space.
  • Our final solution, 4DHands, achieves robust 4D hand mesh recovery in real-world scenarios, serving as a versatile solution to handle hand inputs in various forms, whether the images contain single or two hands, in single frames or sequences.
@@ 2.3 Temporal Hand Mesh Recovery
However, these methods rely solely on sequence data. Our approach introduces a Spatio-Temporal Interaction Reasoning (SIR) module, enabling our method to not only recover accurate two-hand interacting meshes from single-frame input but also produce smooth and stable 4D hand recovery from occluded and motion-blurred temporal input.
@@ 3 Method
Figure 2: Overview of our 4DHands framework. 4DHands is a transformer-based network, which takes monocular images as input and estimates two-hand meshes with their relative positions.
In this Section, we present the technical details of 4DHands. As illustrated in Fig 2, 4DHands is a transformer-based network, which takes hand images as input and estimates the hand meshes, as well as their relative position. In the proposed transformer, the key components include Relation-aware Two-hand Tokenization (RAT) and Spatio-temporal Interaction Reasoning (SIR). As will be discussed shortly, the tokenization and reasoning strategies make 4DHands a generic solution to handle hand images in various forms, regardless of the images with single or two hands, in a single frame or sequences.
@@ 3.1 Problem Formulation
Given a single-frame or sequential image inputs $I$, our goal is to reconstruct the hand meshes and their relative position. We leverage the widely used parametric hand model MANO [41]. MANO defines a function $\mathbf{M}(\theta,\beta)$ that takes two parameters pose $\theta\in\mathbb{R}^{48}$ and shape $\beta\in\mathbb{R}^{10}$ as input, returns 3D hand mesh $V\in\mathbb{R}^{778\times 3}$ and 3D joint keypoints $V\in\mathbb{R}^{21\times 3}$. The outputs of our model are MANO parameters of two hands: $\theta_{R},\theta_{L},\beta_{R},\beta_{L}$, by which we can regress meshes and joints of each hand, and a vector $\upsilon$ to represent the joint root distance of two hands. We define $\upsilon$ to be the location of the left root joint location when we subtract the joints of both hands from the right hand’s 3D root joint location. Note that our method can robustly handle both single-hand and two-hand inputs. When there is only a single hand visible in the images, our method is simply degraded to the estimation of the mesh of one hand.
@@ 3.2 Relation-aware Two-Hand Tokenization
Existing approaches to monocular hand mesh recovery typically take the cropped hand images as input, but neglect the hand relationship in the images. Besides, these approaches handle two-hand and single-hand scenarios in different manners, resulting in jittering artifacts when the two hands operate in interaction and separately. To handle these issues in our transformer-based architecture, we propose a Relation-aware Two-Hand Tokenization (RAT) method to take the relation of two hands into consideration and handle inputs in various scenarios. Following recent state-of-the-art practices in 3D human/hand mesh recovery [51, 15, 39], we utilize ViT [9] as the visual encoder to tokenize image patches. In contrast to these methods, we handle both the single-hand and two-hand inputs in a unified manner and consider embedding positional relation information of each hand into its tokens during the tokenization process.
@@
Specifically, for each input image, we follow ViT [9] to segment it into $H\times W$ image patches and subsequently feed them into the ViT backbone, resulting in a $H\times W\times C$ vision token map. To enrich each token map with relationship information, we further calculate relation maps $\{\mathcal{R}_{R2L},\mathcal{R}_{L2R}\}$ and embed them into the final image tokens, so that they can contain both the patch-wise visual knowledge and patch-to-patch relation information. Basically, the relation maps are the composition of the relative distance map and the overlapping map. Without losing generalization, we describe the right-hand relation map $\mathcal{R}_{R2L}$ in detail as follows. The left-hand relation map $\mathcal{R}_{L2R}$ can be calculated similarly.
@@
where $\oplus$ denotes the concatenation operation. Finally, the relation map acts as the relative positional embedding and enhances the original ViT tokens $F_{R}$ by concatenating them as the relation-aware tokens $F^{+}_{R}$.
3.3 Spatio-temporal Interaction Reasoning
After obtaining the image tokens, we further leverage a Spatio-temporal Interaction Reasoning (SIR) module to infer the hand poses and the relative position between two hands. Specifically, we flatten and concatenate the tokens of two hands to a sequence $M^{*}\in\mathbb{R}^{2HW\times C}$. The token sequence is further fed into attention modules to fuse both the spatial and temporal information.
Adaptive Spatial Fusion
@@ 3.3 Spatio-temporal Interaction Reasoning
To enhance the adaptability of our model, we first employ a self-attention transformer to proceed with the token sequence $M^{*}$. As the relation-aware tokens are embedded with both visual and relation features, the self-attention module can adjust its weight adaptively based on visual and relative positional associativity and mitigate the interdependence of bilateral hand information. For instance, when relation-aware tokens indicate that the two hands are far away from each other, the self-attention can mainly interact with the tokens of each hand individually. The outputs from the self-attention module are further fed into a transformer decoder, resulting in a global feature $G$ corresponding to each hand.
Temporal Fusion
@@
When dealing with the sequential inputs, the global features of each frame are concatenated along the temporal dimension to form the temporal feature $\{G_{t}\}_{t=1}^{T}$, which will be further proceeded by a temporal decoder. By fusing the temporal information, our model can predict the hand poses of intermediate frames more robustly in handling disturbances caused by object occlusions and dynamic blurring. Note that our method is also compatible with single-frame inputs. In this case, we stack the global features repetitively to obtain the feature $\{G_{t}\}$ and feed it into the temporal decoder.
@@
After the feature fusion of both spatial and temporal dimensions, we can obtain the integrated features $G^{*}_{R}$ and $G^{*}_{L}$ of the right and left hand, respectively. Finally, these features are proceeded by an MLP-based regression network to estimate the MANO parameters of two hands and their relative position.

### v3

# OmniHands: Towards Robust 4D Hand Mesh Recovery via A Versatile Transformer

@@ Abstract
In this paper, we introduce OmniHands, a universal approach to recovering interactive hand meshes and their relative movement from monocular or multi-view inputs. Our approach addresses two major limitations of previous methods: lacking a unified solution for handling various hand image inputs and neglecting the positional relationship of two hands within images. To overcome these challenges, we develop a universal architecture with novel tokenization and contextual feature fusion strategies, capable of adapting to a variety of tasks. Specifically, we propose a Relation-aware Two-Hand Tokenization (RAT) method to embed positional relation information into the hand tokens. In this way, our network can handle both single-hand and two-hand inputs and explicitly leverage relative hand positions, facilitating the reconstruction of intricate hand interactions in real-world scenarios. As such tokenization indicates the relative relationship of two hands, it also supports more effective feature fusion. To this end, we further develop a 4D Interaction Reasoning (FIR) module to fuse hand tokens in 4D with attention and decode them into 3D hand meshes and relative temporal movements. The efficacy of our approach is validated on several benchmark datasets. The results on in-the-wild videos and real-world scenarios demonstrate the superior performances of our approach for interactive hand reconstruction. More video results can be found on the project page: https://OmniHand.github.io.
Figure 1: The proposed method, OmniHands, can robustly recover interactive hand meshes and their relative movement from monocular inputs.
@@ 1 Introduction
Hand mesh recovery is a fundamental technology for numerous advanced applications like virtual/augmented reality, human-computer interaction, robotics, and embodied AI.
Despite the remarkable progress achieved by existing methods [32, 34, 64] in this field, three major challenges remain unresolved. First, there is a lack of a universal method that can handle various input statuses for both single-handed and two-handed input. When using existing single-hand reconstruction models to reconstruct two hands, while they can individually reconstruct each hand, they fail to address the issue of hand occlusion and cannot compute the relative distance between the hands. On the other hand, current two-hand reconstruction models, due to their use of single images containing both hands as input, are limited to reconstructing scenarios where the hands are in close proximity, thereby exhibiting poor generalizability. Second, current hand recovery methods address specific subproblems such as monocular hand reconstruction, video-based reconstruction, and multi-view reconstruction, however, a unified and general model to handle these tasks is still lacking. Third, although current hand reconstruction methods perform well on test data, their effectiveness in in-the-wild scenarios is limited due to the inadequate size of training datasets, resulting in poor practical value.
The success of recent approaches in transformer-based large-scale models such as GPT-4[1], ViTpose [61], HaMeR [45] provided us with the insight to investigate the usage of simple and high-capacity models along with large-scale in-the-wild data. Therefore, we proposed OmniHands, a transformer-based universal solution to recover interactive hand meshes cross omniform scenarios. The transformer possesses versatility, enabling it to perform unified forward operations on different input forms such as single-frame input, temporal input, and multi-view input. Additionally, multi-layer transformers’ ability to process large datasets allows us to train a generalized model that remains robust in real-world scenarios across various lighting conditions, backgrounds, and interaction poses.
In the proposed method OmniHands, we unified the following aspects to achieve robust 4D hand mesh recovery: i) Status of Hand Images. The input can be single hand images or interactive hands images. In previous inter-hand recovery methods, a single image containing both hands is typically used as input. However, to accommodate both single-hand and inter-hand recovery scenarios, we need to unify the input format that handles both types of input while preserving the relative positions of the hands. ii) View Points. Inputting sequences of images from multiple camera perspectives captured at the same moment to enhance details through information supplementation across different viewpoints. iii) Temporal Sequences. Inputting a monocular dynamic sequence, make more accurate predictions by evaluating hand movements under conditions of motion blur and object occlusion. Since the multi-view data and temporal data can both be seen as contextual sequences, a unified module can be applied to fuse the contextual information.
To address these issues, we proposed a Relation-aware Two-Hand Tokenization (RAT) method and a 4D Interaction Reasoning (FIR) module to unifiedly process features in spatial context or temporal context.
@@
In Relation-aware Two-Hand Tokenization (RAT), we crop and tokenize the RGB information of each hand individually, while maintaining hand-hand relative position information in tokens. Such a tokenization method brings the following benefits threefold. i) RAT is not reliant on the co-occurrence of two hands and can thus handle both single-hand and two-hand inputs in a unified manner. ii) RAT takes the overlapped status of two hands into account and adjusts the tokens accordingly, which allows the network to learn a more concentrated distribution of the hand tokens regardless of the various scales and positions. iii) RAT explicitly encodes the relative positions of two hands into the tokens, facilitating the reconstruction of close and intricate interaction in real-world scenarios.
Upon obtaining the image tokens, we integrate a Context Interaction Reasoning (FIR) module to unifiedly fuse contextual relations within the context channel. These contextual relations could be temporal (over time) or spatial (from multiple camera views). Thanks to the proposed tokenization method RAT, FIR can fuse both the contextual information more effectively as the hand tokens indicate the relative relationship of two hands, thereby enhancing the robustness of hand mesh and contextual movement recovery.
@@
  • We propose a 4D Interaction Reasoning (FIR) module to fuse spatial or temporal hand features. The collaboration of RAT and FIR further enhances the stability of hand mesh and temporal movement recovery.
  • Our final solution, OmniHands, achieves robust hand mesh recovery in real-world scenarios, serving as a versatile solution to handle hand inputs in various forms, whether the images contain single or two hands, in single frames or temporal sequences or multi-view sequences.
@@ 2.3 Temporal Hand Mesh Recovery
However, these methods rely solely on sequence data. Our approach introduces a 4D Interaction Reasoning (FIR) module, enabling our method to not only recover accurate two-hand interacting meshes from single-frame input but also produce smooth and stable 4D hand recovery from occluded and motion-blurred temporal input.
2.4 Multi-View Hand Mesh Recovery
Recovering hands from multi-view input is a similar task as multi-view human reconstruction. Many recent researches have been proposed for multi-view 3D human pose estimation  [46, 11, 53, 44, 72]. Some of the methods [46, 11, 53] focus on matching the projection of 3D key-points with their corresponding 2D key-points from multiple camera views, and some [44, 72] focus on fusing the 2D visual features to build a 3D structure.
Recent research on multi-view hand reconstruction has seen various innovative approaches. Some researchers [10, 20] utilize implicit fields to recover the shape and appearance of hands. Yang et al. [63] employ hand vertices as a point cloud embedded in the intersection volume of frustums from multiple cameras to integrate multi-view information. Wang et al. [58] adapt multi-view hand data to enhance the performance of single-view hand reconstruction. Additionally, Gan et al. [17] propose a neural rendering scheme to refine both the hand mesh and textures. However, most multi-view methods require extrinsic and intrinsic calibration information for each camera to accurately determine the 3D positions of hands. In contrast, our approach is capable of utilizing only visual information to recover precise hand meshes.
@@ 3 Method
Figure 2: Overview of our OmniHands framework. OmniHands is a transformer-based network, which takes various forms of inputs and estimates two-hand meshes with their relative positions. General modules are used to process different forms of input data. For simplicity of representation, since the hand tokens of both hands will undergo the same processing in FIR, the hand tokens $G^{*},L^{*}$ in the diagram represent those of one hand. In the top left corner we present our training set, indicating the data volume and whether including Interaction, Time sequences, Multi-view, Real-World data.
In this Section, we present the technical details of OmniHands. As illustrated in Figure 2, OmniHands is a transformer-based network, which can accept multiple forms of input: single-frame images, videos, and multi-view images to estimates the hand meshes, as well as their relative position. In the proposed transformer, the key components include Relation-aware Two-hand Tokenization (RAT) and 4D Interaction Reasoning (FIR). As will be discussed shortly, the tokenization and reasoning strategies make OmniHands a generic solution to handle hand images in various forms, regardless of the images with single or two hands, in a single frame or temproal sequences or multi-view sequences.
@@ 3.1 Problem Formulation
Given a single-frame or a temporal sequence or a multi-view sequence of images $I$, our goal is to reconstruct the hand meshes and their relative position. We leverage the widely used parametric hand model MANO [48]. MANO defines a function $\mathbf{Mn}(\theta,\beta)$ that takes two parameters pose $\theta\in\mathbb{R}^{48}$ and shape $\beta\in\mathbb{R}^{10}$ as input, returns 3D hand mesh $V\in\mathbb{R}^{778\times 3}$ and 3D joint keypoints $V\in\mathbb{R}^{21\times 3}$. The outputs of our model are MANO parameters of two hands: $\theta_{R},\theta_{L},\beta_{R},\beta_{L}$, by which we can regress meshes and joints of each hand, and a vector $\upsilon$ to represent the joint root distance of two hands. We define $\upsilon$ to be the location of the left root joint location when we subtract the joints of both hands from the right hand’s 3D root joint location. Note that our method can robustly handle both single-hand and two-hand inputs. When there is only a single hand visible in the images, our method is simply degraded to the estimation of the mesh of one hand.
@@ 3.2 Relation-aware Two-Hand Tokenization
Existing approaches to hand mesh recovery typically take the cropped hand images as input, but neglect the hand relationship in the images. Besides, these approaches handle two-hand and single-hand scenarios in different manners, resulting in jittering artifacts when the two hands operate in interaction and separately. To handle these issues in our transformer-based architecture, we propose a Relation-aware Two-Hand Tokenization (RAT) method to take the relation of two hands into consideration and handle inputs in various scenarios. Following recent state-of-the-art practices in 3D human/hand mesh recovery [60, 19, 45], we utilize ViT [12] as the visual encoder to tokenize image patches. In contrast to these methods, we handle both the single-hand and two-hand inputs in a unified manner and consider embedding positional relation information of each hand into its tokens during the tokenization process.
@@
Specifically, for each input image, we follow ViT [12] to segment it into $H\times W$ image patches and subsequently feed them into the ViT backbone, resulting in a vision token map $M\in\mathbb{R}^{H\times W\times C}$ To enrich each token map with relationship information, we further calculate relation maps $\{\mathcal{R}_{R2L},\mathcal{R}_{L2R}\}$ and embed them into the final image tokens, so that they can contain both the patch-wise visual knowledge and patch-to-patch relation information. Basically, the relation maps are the composition of the relative distance map and the overlapping map. Without losing generalization, we describe the right-hand relation map $\mathcal{R}_{R2L}$ in detail as follows. The left-hand relation map $\mathcal{R}_{L2R}$ can be calculated similarly.
@@
where $\oplus$ denotes the concatenation operation. Finally, the relation map acts as the relative positional embedding and enhances the original ViT tokens $M_{R}$ by concatenating them as the relation-aware tokens $\mathcal{M}_{R}\in\mathbb{R}^{HW\times C}$.
$\mathcal{M_{R}}=Flatten(M_{R}+\mathcal{R}_{R2L})$ (5)
Cross-hand Feature Fusion
With the token maps, we can fuse the cross-hand features by a Bimanual Decoder. Specifically, for each single frame we concatenate the tokens of two hands to $\mathcal{M}=\mathcal{M_{R}}\oplus\mathcal{M_{L}}$. Then $\mathcal{M}$ is further fed into attention modules to fuse inter-hand information. To enhance the adaptability of our model, we first employ Spatio Encoder to proceed with the token sequence $\mathcal{M}$:
$\mathcal{M^{*}}=FFN(Softmax(\frac{Q(\mathcal{M})K(\mathcal{M})}{\sqrt{dim(\mathcal{M})})})V(\mathcal{M}))$ (6)
where $Q,K,V$ respectively represent the left-multiplied query, key, and value matrices.
@@
As the relation-aware tokens are embedded with both visual and relation features, the self-attention module can adjust its weight adaptively based on visual and relative positional associativity and mitigate the interdependence of bilateral hand information. For instance, when relation-aware tokens indicate that the two hands are far away from each other, the self-attention can mainly interact with the tokens of each hand individually. The outputs $\mathcal{M}^{*}$ from the self-attention module are further fed into Spatio Decoder, where we take a constant vector $\alpha\in\mathbb{R}^{C}$ as query vector, resulting in a global hand token $G\in\mathbb{R}^{C}$ corresponding to each hand:
$G=FFN(Softmax(\frac{Q(\alpha)K(\mathcal{M}^{*})}{\sqrt{dim(\mathcal{M}^{*})})})V(\mathcal{M}^{*}))$ (7)
Then the hand tokens are taken for further feature fusion according to the specific task demands.
3.3 4D Interaction Reasoning
Regardless of whether the input consists of monocular temporal frames or multi-view non-temporal frames, we can consider the set of inputs as contextually relative, therefore employ similar methods to correlate the information contained within them. Consequently, we aim to utilize a unified module 4D Interaction Reasoning (FIR) module in order to fuse the contextual knowledge to infer the hand poses and the relative position between two hands. We will discuss the processing differences specific to temporal inputs and multi-view inputs.
Temporal Context Fusion
When dealing with the temporal sequential inputs, we use $\{G_{t}\}_{t=1}^{T}$ to represent a temporal sequence of one of the two hands, for simplicity, and $\{G_{t}\}_{t=1}^{T}$ will be further proceeded by Global Context Decoder.
@@
By fusing the sequential information, our model can predict the hand poses of intermediate frames more robustly in handling disturbances caused by object occlusions and dynamic blurring. Note that our method is also compatible with single-frame inputs. In this case, we stack the hand tokens repetitively to obtain the sequential token $\mathcal{G}\in\mathbb{R}^{T\times C}$. Since the temporal input has a sequential order, we need to add position embedding to $\mathcal{G}$ before inputting it into Collective Decoder to get global feature. We only output the results of the middle $\frac{t}{2}$-th frame because, when predicting the middle frame, both past and future information is available, leading to more accurate predictions:
$G^{*}=FFN(Softmax(\frac{Q(\alpha)K(\mathcal{G})}{\sqrt{dim(\mathcal{G})})})V(\mathcal{G})$ (8)
@@
After the feature fusion of both spatial and temporal dimensions, we can obtain the integrated features $G^{*}_{R},G^{*}_{L}\in\mathbb{R}^{C}$ as right hand collective token and left hand collective token, respectively. Finally, these features are proceeded by a Global Regressor, an MLP-based regression network to estimate the MANO parameters $\theta_{R},\theta_{L},\beta_{R},\beta_{L}$ of two hands and their relative position $\upsilon$, and camera parameters $\pi$ for 2D projection.
Multi-view Context Fusion
When inputting frames of multiple camera views, we get hand tokens $\mathcal{G}\in\mathbb{R}^{T\times C}$ in the same way as mentioned above, where $T$ is the number of cameras instead of time length. Then concatenating them to obtain the sequential global feature $\mathcal{G}$. Different with Temporal Context Fusion, since the camera parameters are not required as input, there isn’t a fixed order of input frames and can be combined in any arbitrary sequence, therefore we don’t add position embedding to $\mathcal{G}$ in this case. In temporal context fusion, we utilize information from preceding and succeeding frames to predict an accurate result of the intermediate frame. However, in multi-view context fusion, since there is no sequential order, the distance between frames is equivalent for all, enabling the mutual supplementation of information from each frame. Consequently, we can output predictions for all frames, with the collective information contributing to a precise result:
$\mathcal{L}^{*}=FFN(Softmax(\frac{Q(\mathcal{G})K(\mathcal{G})}{\sqrt{dim(\mathcal{G})})})V(\mathcal{G})$ (9)
We get individual tokens $\mathcal{L}^{*}\in\mathbb{R}^{T\times C}$ from Individual Encoder (Eq 9) and collective features $G^{*}\in\mathbb{R}^{\times C}$ from Collective Decoder (Eq 8) after fusion. Given that all frames are taken at the same time from different views, the gesture pose $\theta_{3:45}$ and shape $\beta$ should be the same, while the root orientation $\theta_{0:3}$, camera parameters $\pi$ and root offset $\upsilon$ varying. So we feed global feature $G^{*}$ into Collective Regressor, to regress the consistent parameters $\theta_{3:45}$ and shape $\beta$, and $\mathcal{L}^{*}$ into MLP-based Regressor to regress the inconsistent parameters $\theta_{0:3}$, $\pi$, $\upsilon$ for each frame.
Although we predict the root orientations and offset vectors for each frame individually, the relative orientation and distance between the right hand and left hand should remain invariant across all frames. This means that when we transform all right hands to the zero position and zero rotation, the positions and orientations of the left hands should remain consistent. Therefore, we apply a constraint (Eq 17) to maintain this invariance.